OpenStudy (anonymous):

Which of the following lines is perpendicular to y = –5x – 2?

OpenStudy (anonymous):

its slope=-5 slope of the line perpendicular to above line is-1/-5=1/5 equation of any line perpendicular to given line is y=1/5 x+c,or 5y=x+5c,5y=x+C or y-a=1/5 (x-b),where (a,b) is the point through which the required line passes. 5y-5a=x-b or 5y=x+5a-b or 5y=x+C